OpenStudy (anonymous):

solve. 4/z+6 = 3/z-7

OpenStudy (anonymous):

This is a proportion. Find the cross products. 4(z-7)=3(z-7). Then distribute. Solve for z.

OpenStudy (anonymous):

4(z-7)=3(z-6) sorry.

OpenStudy (anonymous):

It is a rational expression

OpenStudy (anonymous):

Right, to remove the the denominators we can treat it as a proportion.

OpenStudy (anonymous):

the math is the same as solving something like 2/x = 9/14 9x = 2*14

OpenStudy (anonymous):

\[{4 \over z+6} = {3 \over z-7}\] 4(z-7) =3(z+6) 4z-28=3z+18 and so on

OpenStudy (anonymous):

so is the answer -10/7 ? I'm not sure if I understand it correctly

OpenStudy (anonymous):

subtract 3z from both sides 1z -28 = 18 add 28 to both sides 1z = 46

OpenStudy (anonymous):

Check out this example http://www.purplemath.com/modules/solvrtnl.htm